OpenStudy (anonymous):

A bag contains 1 blue, 2 white, and 2 red marbles. Match the columns. Drag the letter after a compound event in the left column and drop it before the matching probability in the right column. Select Submit to see if you are correct. 1- P(red and red), with replacement 2- P(white and white), without replacement 3- P(blue and red), with replacement 4- P(white and red), without replacement A- 16% B- 8% C- 20% D- 10% (Know that you cannot have the same answer for more than one match. Example: You CAN'T have: 1-A, 2-A, 3-C, 4-D)

OpenStudy (anonymous):

@HawkCrimson @Mertsj

OpenStudy (anonymous):

...it's so long.. i dont like doing long questions..

OpenStudy (anonymous):

summerize it :P

OpenStudy (anonymous):

ah @Mertsj will take over :P

OpenStudy (mertsj):

@ImABibliophile How many marbles are in the bag?

OpenStudy (anonymous):

5

OpenStudy (mertsj):

How many of them are red?

OpenStudy (anonymous):

2

OpenStudy (mertsj):

You reach in the bag and draw out a marble. What is the probability it is red?

OpenStudy (anonymous):

40% chance... i think.

OpenStudy (mertsj):

Well you could make like simple and just say 2/5

OpenStudy (anonymous):

Ya, the answers are in percents so Im practicing :D

OpenStudy (mertsj):

Ok. You put the red marble back. Now how many marbles are in the bag?

OpenStudy (anonymous):

5 since you put it back

OpenStudy (mertsj):

How many of them are red?

OpenStudy (anonymous):

2 out of 5

OpenStudy (mertsj):

So yes, the probability that you draw another red marble is exactly the same as at the beginning since the conditions are the same. So the probability of getting a red marble on the first draw is 2/5. The probability of getting a red marble on the second draw is 2/5. Multiply those individual probabilities. What do you get?

OpenStudy (mertsj):

\[\frac{2}{5}\times \frac{2}{5}=\]

OpenStudy (anonymous):

4/25

OpenStudy (anonymous):

sorry, i take a long time doing problems

OpenStudy (mertsj):

Divide 4 by 25 and change it to a percent.

OpenStudy (anonymous):

so that would be 16%

OpenStudy (anonymous):

and so 1- C

OpenStudy (anonymous):

but now we have without replacement which I am even more confused with

OpenStudy (mertsj):

Ok. We'll start over. There are 5 marbles in the bag, 2 of them are white. What is the probability you will draw out a white marble?

OpenStudy (anonymous):

it says im typing but im not. well... now i am but not before

OpenStudy (anonymous):

40% like the red marbles

OpenStudy (anonymous):

because of there being also two whites like the two reds. so 40% or 2/5

OpenStudy (mertsj):

Well finally, I got back on here. It is very slow, however.

OpenStudy (mertsj):

Yes. 2/5 is the probability of drawing out a white marble. Now you have taken the white marble out of the bag. You did NOT put it back. How many marbles are in the bag now?

OpenStudy (anonymous):

4 marbles are left in the bag

OpenStudy (mertsj):

How many of them are white?

OpenStudy (anonymous):

Since you took out one, there is only 1 left so 1/4 or 25%

OpenStudy (mertsj):

So, again, multiply the individual probabilities: \[\frac{2}{5}\times\frac{1}{4}=\]

OpenStudy (anonymous):

1/10 or 10%

OpenStudy (mertsj):

Good. Do you understand? Can you try the next one?

OpenStudy (anonymous):

I get it. Let me try it. I'll reply the answer back on here so you can check it. :D

OpenStudy (anonymous):

Wait, when they say red, does it mean just one red marble or both?

OpenStudy (mertsj):

Number 3 says to find the probability of drawing a blue marble and a red marble if you put the blue marble back. How many marbles are in the bag? Answer 5 How many marbles are blue? Answer 1 What is the probability of drawing a blue marble? Answer 1/5 Now you put the blue marble back in the bag. How many marbles are in the back? Answer: 5 How many marbles are red? Answer: 2 What is the probability of drawing a red marble? Answer 2/5 What is the product of 1/5 and 2/5? Answer : 2/25 What is 2/25 expressed as a percent? Answer: 8%

OpenStudy (anonymous):

@Mertsj Ya, i tried sending that but my computer froze so at least I got it right!

OpenStudy (mertsj):

Very good. Now see if you can verify that the answer to the fourth one is the remaining choice.

OpenStudy (anonymous):

Okay, so if we take out the reds and whites, we have 4/5. And we don't put them back in. So now we have 2/5 reds and 2/5 whites and 1/1 blue left in the bag. Now here is where I'm confused. I don't know what to do.

OpenStudy (anonymous):

I might understand it now

OpenStudy (mertsj):

Number 4: Probability of white and red and you don't put the white one back. What is the probability of drawing a white marble? 2/5 Since you didn't put it back there are 4 marbles in the bag and 2 of them are red. what is the probability of drawing a red marble: 2/4 = 1/2 Multiply 1/2 and 2/5=2/10=20%

OpenStudy (anonymous):

Ohhhhh... I get it now!

OpenStudy (mertsj):

Very good.

OpenStudy (anonymous):

Thx so much for your help!

OpenStudy (mertsj):

yw